A coil of resistance 100 Ω is connected across a battery of emf 6.0 V. Assume that the heat developed in the coil is used to raise it temperature. If the heat capacity of the coil is 4.0 J/K, how long will it take to raise the temperature of the coil by 15°C?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. R = 100 Ω, Heat capacity of the coil = 4 J/k ∆T = 15°c Heat liberate ⇒ E^2/RT = 4 J/K * 15 ⇒ 6 * 6/100 *t = 60 ⇒ t = 166.67 sec = 2.8 min